OpenStudy (anonymous):

integrate 5x cube minus sin3x

OpenStudy (anonymous):

\[\int\limits ((5x)^3-\sin(3x))dx\] you mean that?

OpenStudy (anonymous):

well if that's what you mean you can make 2 integrals out of it and just integrate the left part and then the right part of it. need more help?

OpenStudy (anonymous):

Yes I do

OpenStudy (anonymous):

well now we have \[\int\limits (5x)^3 dx - \int\limits \sin(3x) dx\] well, ... i could write it out now for you but maybe you can remember some rules about integrating and that will help you. i can calculate it for you but that would not help you much i guess.

OpenStudy (anonymous):

\[5 \int\limits x^3 dx - \int\limits \sin(3x) dx\] now just think of what equation you have to build the derivate of to get x^3 and sin 3x

OpenStudy (anonymous):

x raised to power four over 4

OpenStudy (anonymous):

right

OpenStudy (anonymous):

that's right so we have the left side with \[\frac{125x^4}{ 4 }\] (i forgot the 5^3 before) and what about the right side? what is the derivate of sinus?

OpenStudy (anonymous):

am going to guess here

OpenStudy (anonymous):

minus cos

OpenStudy (anonymous):

sinus --> cosinus --> minus sinus --> minus cosinus --> sinus ( the loop begins anew) so you have sin(3x) that means before that you hade a minus cosinus. therefore what is the solution for the right side?

OpenStudy (anonymous):

plus cosx

OpenStudy (anonymous):

plus cos3x

OpenStudy (anonymous):

good but there is something missing. what happens if you build the derivate of the equation you just wrote? you get:\[-3\sin(3x)\] so you need to divide it with 3 and the whole endresult is: \[\int\limits ((5x)^3-\sin(3x) ) dx = \frac{125x^4} {4} + \frac{\cos(3x)}{3}\]

OpenStudy (anonymous):

why divide by 3

OpenStudy (anonymous):

was 3x diffrentiated

OpenStudy (aravindg):

actually we divided cos 3x by 3 because while taking differential we get multiplied by 3

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Chain_rule that explains the /3

OpenStudy (aravindg):

simply differentiate cos 3x/3 you will get back -sin 3x

OpenStudy (anonymous):

chain rule

OpenStudy (anonymous):

indeed

OpenStudy (anonymous):

so integrate cos3x and divide by diffrention of 3x

OpenStudy (anonymous):

is that it

OpenStudy (anonymous):

well yes, but i wouldn't call it like that. just: but it's fine :)

OpenStudy (anonymous):

Thank you